# Disentangling Rich Dynamics from Feature Learning: A Framework for Independent Measurements

### v2

@@ D.3 Generality of our metric beyond neural collapse
As discussed in the main text, our metrics are defined in terms of functions (random variables) while neural collapse is defined by the feature vectors of the training set. Working in the function space, we can measure the quality Eq. 5, which is inaccessible from formalism using training feature vectors.
@@
Another important difference is that we focus on the learned function instead of mean class label vectors. This allows more independence beyond neural collapse, allowing the framework to be applied on broader set of tasks: tasks with orthogonal and isotropic target functions form a broader class than balanced classification tasks. Regression task with scalar output is an example.
@@
In Fig. 12, we train ResNet18 on MNIST as a regression task with scalar output ($f^{*}:\mathcal{X}\rightarrow\mathbb{R}$), where a digit $i$ is correctly predicted if $i-0.5<f(x)<i+0.5$. Unlike $C$-way classification, which typically uses $C$ features, the regression model reaches the rich regime with just two — one of which is constant. Neural collapse formalism, focusing on class labels, does not readily extend to regression tasks, demonstrating that our function-based formalism is more general, label-independent, and applicable beyond classification.
@@
Figure 12: Rich dynamics for regression problem with scalar output. ResNet18 was trained on MNIST via regression (blue) and classification (orange). For regression, the target/learned functions are scalar output functions, and an image $x$ with digit $i$ is correctly classified if $i-0.5<f(x)<i+0.5$. It can be seen that both models are in the rich regime, using the minimum number of features.
@@
Finally, the independence from the class labels allows another benefit of being applicable during training. Neural collapse is defined only during the terminal stage of training and assumes 100% classification. A measure without perfect classification can lead to unstable values, typically rising from the pseudo-inverse of $\Sigma_{b}$ (See also Appendix F). The NC1 measure shows a dramatic change in logarithmic scale for experiments in Fig. 2(a,b) while our measure is more robust to such differences.
Appendix E Examples of our visualization framework
@@
Rich dynamics $\neq$ generalization.
In Fig. 1, we showed an example where rich dynamics led to poorer performance. Fig. 13 shows a similar experiment in which both rich and lazy models achieve near-perfect performance ($>95\%$ test accuracy). Our visualization clearly shows the different usages of features even when the performances are similar.
@@
Table 1 shows that low-rank inductive bias is strong in vision tasks, where the number of significant features always equals $C$ (the number of orthogonal functions). See Eq. 9 in Section B.4 for the $D_{eff}(\rho)$ expression. We also observe that our measure $\mathcal{D}_{LR}$ is small in these setups.
@@
Fig. 14 shows that weight decay indeed aids low-rank representation and smaller $\mathcal{D}_{LR}$. However, the model without any weight decay already shows sufficiently low-rank representations, suggesting that dynamical inductive bias, not the weight decay is the main driving cause of low-rank representations.
@@
Figure 15: Effects of learning rates on feature learning and generalization. We explore the effect of changing the learning rate for a ResNet18 on CIFAR100. Note that a lower learning rate leads to a lazier regime and poorer performance. The best performance is for the intermediate learning rate in the rich regime with the tightest $\mathcal{D}_{LR}$ measure.
Effect of learning rate on rich dynamics.
Fig. 15 shows that a smaller learning rate induces lazier dynamics, while a too large learning rate (largest without gradient blow up) also deviates from the rich regime. The optimal learning rate correlates with the best performance and the richest dynamics. However, no extensive searches were performed on learning rates.
@@
As studied in Saxe et al. [55], the architecture and dataset pair can influence the performance even when dynamics are greedy (race toward shared representation). In Fig. 16, we examine the dynamics of ResNet18 and a 4-layer MLP trained on CIFAR-10. In Fig. 16(a), the training dynamics is restricted to the first 10 significant features, consistent with theoretical work in linear neural networks. In Fig. 16(b), the training is mainly focused on the first 10 features, but we observe that more features are used as training progresses, more similar to the dynamics of linear models. We speculate that the lack of ability to feature learn in the earlier layers leads to the use of additional features and thus leads to lazier training dynamics.
@@
Fig. 17 shows that our method extends to models trained on cross-entropy loss, while the mathematical interpretation is less straightforward.
Figure 17: Training with cross-entropy loss. We apply our visualization method to a ResNet18 trained on CIFAR-10 with cross-entropy loss. Similar to the MSE-trained model, it achieves small value of $\mathcal{D}_{LR}$ and shows concentration of feature usage and significance. Insets display all 512 features rather than just the top 25.
Appendix F Empirical details
Here, we share information on our practical setups and their statistical significance.
F.1 Statistical significance
@@
Train acc. $\uparrow$ 100 (±$(\pm 0)$98.7 (±$(\pm 0.11)$
Test acc. $\uparrow$ 10.0 (±$(\pm 4.3\cdot 10^{-2})$74.4 (±$(\pm 0.19)$
$\mathcal{D}_{LR}$downarrow $8.7\cdot 10^{-3}$ (±$(\pm 4.0\cdot 10^{-4})$ $6.3\cdot 10^{-1}$ (±$(\pm 3.2\cdot 10^{-3})$
Table 3: Statistical significance of Figure 2(a)
@@
Train Acc.$\uparrow$ 10.0 (±$(\pm 0.8)$10.2(±$(\pm 0.27)$%
Test Acc.$\uparrow$ 9.85 (±$(\pm 1.0)$10.0(±$(\pm 0.33)$%
$\bm{\mathcal{D}_{LR}}$$\downarrow$ 0.59 (±$(\pm 1.6\cdot 10^{-2})$ 1.0 (±$(\pm 6.4\cdot 10^{-5})$
$\mathcal{S}_{init}$$\downarrow$ 1.0 (±$(\pm 0)$ 0.20 (±$(\pm 1.1\cdot 10^{-2})$
$\|\theta\|_{F}^{2}$$\downarrow$ 3.1 $\cdot 10^{3}$ (±$(\pm 3.5\cdot 10^{2})$ 2.2 $\cdot 10^{-5}$ (±$(\pm 2.5\cdot 10^{-9})$
$NC1$$\downarrow$ $1.2\cdot 10^{5}$ (±$(\pm 7.2\cdot 10^{2})$ $7.5\cdot 10^{-14}$ (±$(\pm 1.5\cdot 10^{-14})$
$Tr(\Sigma_{W})$$\downarrow$ $15.6$ (±$(\pm 1.4)$ $7.1\cdot 10^{-18}$ (±$(\pm 1.1\cdot 10^{-18})$
Table 4: Statistical significance of Figure 2(b)
@@
Train Acc.$\uparrow$ 100 (±$(\pm 0)$100 (±$(\pm 0)$100 (±$(\pm 0)$%
Test Acc.$\uparrow$ 92.7 (±$(\pm 3.4\cdot 10^{-2})$92.4 (±$(\pm 5.2\cdot 10^{-2})$88.3 (±$(\pm 3.7\cdot 10^{-1})$%
$\bm{\mathcal{D}_{LR}}$$\downarrow$ $\mathbf{4.9\cdot 10^{-2}}$ (±$(\pm 4.4\cdot 10^{-3})$ $\mathbf{1.1\cdot 10^{-1}}$ (±$(\pm 9.4\cdot 10^{-3})$ $\mathbf{5.6\cdot 10^{-1}}$ (±$(\pm 6.1\cdot 10^{-2})$
$\mathcal{S}_{init}$$\downarrow$ $6.8\cdot 10^{-2}$ (±$(\pm 2.1\cdot 10^{-3})$ $4.1\cdot 10^{-2}$ (±$(\pm 4.5\cdot 10^{-3})$ $5.2\cdot 10^{-2}$ (±$(\pm 1.8\cdot 10^{-2})$
$\|\theta\|_{F}^{2}$$\downarrow$ $3.4\cdot 10^{3}$ (±$(\pm 1.1\cdot 10^{1})$ $3.2\cdot 10^{3}$ (±$(\pm 4.4)$ $3.2\cdot 10^{3}$ (±$(\pm 2.0)$
$NC1$$\downarrow$ $2.3\cdot 10^{4}$ (±$(\pm 4.1\cdot 10^{3})$ $3.2\cdot 10^{3}$ (±$(\pm 4.3\cdot 10^{2})$ $8.1\cdot 10^{2}$ (±$(\pm 1.4\cdot 10^{2})$
$Tr(\Sigma_{W})$$\downarrow$ $2.0$ (±$(\pm 0.25)$ $3.1\cdot 10^{-1}$ (±$(\pm 2.3\cdot 10^{-2})$ $1.2\cdot 10^{-1}$ (±$(\pm 1.2\cdot 10^{-2})$
F.2 Discussion on NC1 measure
In Tables 3 and 4, the NC1 measure is numerically unstable due to the pseudo-inverse of $\Sigma_{b}$. For stability, We used $\Sigma_{b}^{\dagger}\approx(\Sigma_{b}+10^{-4}I)^{\dagger}$. In addition to NC1, we consider a similar measure $Tr(\Sigma_{W})$ [6] for collapse of within class variability, which is also reported in Tables 3 and 4. The measure $Tr(\Sigma_{W})$ also shows a similar trend to NC1.
F.3 Dataset details
We use publicly available datasets, including MNIST [56] and CIFAR-10/100 [57] from Pytorch [58], and mod $p$ division task from https://github.com/teddykoker/grokking.
F.4 Model details
Our model implementations are based on publicly available code assets, including VGG16 from PyTorch [58], ResNet18 from Nakkiran et al. [59], and the modular division task from https://github.com/teddykoker/grokking.
As described in the main text, we only trained ours models on MSE loss where the targets are one-hot vectors (up to scaling constant). The constant $\alpha$ is used to scale the output $y\rightarrow y/\alpha$.

### v3

@@ D.3 Generality of our metric beyond neural collapse
As discussed in the main text, our metrics are defined in terms of functions (random variables) while neural collapse is defined by the feature vectors of the training set. Working in the function space, we can measure the quality Eq.˜5, which is inaccessible from formalism using training feature vectors.
@@
Another important difference is that we focus on the learned function instead of mean class label vectors. This allows more independence beyond neural collapse, allowing the method to be applied on broader set of tasks: tasks with orthogonal and isotropic target functions form a broader class than balanced classification tasks. Regression task with scalar output is an example.
@@
In Fig.˜10, we train ResNet18 on MNIST as a regression task with scalar output ($f^{*}:\mathcal{X}\rightarrow\mathbb{R}$), where a digit $i$ is correctly predicted if i0.5<f(x)<i+0.5i-0.5<f(x)<i+0.5. Unlike $C$-way classification, which typically uses $C$ features, the regression model reaches the rich regime with just two — one of which is constant. Neural collapse formalism, focusing on class labels, does not readily extend to regression tasks, demonstrating that our function-based formalism is more general, label-independent, and applicable beyond classification.
@@
Figure 10: Rich dynamics for regression problem with scalar output. ResNet18 was trained on MNIST via regression (blue) and classification (orange). For regression, the target/learned functions are scalar output functions, and an image $x$ with digit $i$ is correctly classified if i0.5<f(x)<i+0.5i-0.5<f(x)<i+0.5. It can be seen that both models are in the rich regime, using the minimum number of features.
@@
Finally, the independence from the class labels allows another benefit of being applicable during training. Neural collapse is defined only during the terminal stage of training and assumes 100% classification. A measure without perfect classification can lead to unstable values, typically rising from the pseudo-inverse of $\Sigma_{b}$ (See also Appendix˜H). The NC1 measure shows a dramatic change in logarithmic scale for experiments in Fig.˜11(a,b) while our measure is more robust to such differences.
D.4 Isotropicity assumption
As stated at the beginning of Section 2 and in the limitation paragraph, our paper’s scope is on isotropic outputs. This is a slightly weaker assumption compared to prevailing standards in neural collapse literature (Papyan et al., 2020; Mixon et al., 2020; Han et al., 2022; Jacot et al., 2025), as discussed above. Because simplex ETF structures rely significantly on the symmetry, proposing a plausible extension for imbalanced tasks is an active area of research on its own (Fang et al., 2021; Yang et al., 2022; Yan et al., 2024; Gao et al., 2024). For example, imbalanced targets suffer from the minority collapse where multiple classes collapse to a single vector for a finite dataset (Fang et al., 2021). This becomes particularly a problem for a long-tailed target (e.g., natural language dataset), which is also an active area of research (Gao et al., 2024).
As another challenge, the eigenvalue distribution of features now depends on many factors such as weight decay. Assume a diagonal last layer with $p=C$. Assuming anisotropic target $f^{*}_{k}(x)=a_{k}e_{k}(x)$, the intensity $a_{k}$ can either come from the $\|\phi_{k}(x)\|$ or the last layer weight $w_{k}$. In other words,
$a_{k}=w_{k}\|\phi_{k}(x)\|.$ (57)
For strong weight decay, the weights $w_{1},\ldots,w_{p}]$ become more isotropic and anisotropy are absorbed in $\|\phi_{k}(x)\|$. For two-layer diagonal linear networks in rich regime with isotropic input ($x_{k}\sim\mathcal{N}(0,1)$), $w_{k}\propto|\phi_{k}(x)|$ where anisotropy is distributed evenly between the weights and features. Since our metric compares $\phi(x)$ to $f(x)$, these differences, though yielding the same output, returns different levels of dynamical richness. But then, what representation shapes are favored by rich dynamics? While symmetry from an isotropic target and normalized CKA have let us bypass this issue, introducing anisotropy and factors such as weight decay complicates the dynamics and requires more concise understanding of the rich dynamics.
Appendix E Calculating the metric and visualization in practice
E.1 Metric
In the main text, we used function and operator formalism. In practice, we must use finite approximation. We can use the CKA equation Eq.˜7
$CKA(A,B)=\frac{\mathrm{Tr}(c(A)c(B))}{\|c(A)\|_{F}\|c(B)\|_{F}},$ (58)
but use $n\times n$ feature matrices for $A$ and $B$ instead of operators. Let $A^{1/2}\in\mathbb{R}^{p\times n}$ be feature vectors (of the penultimate layer) of $n$ samples and $B^{1/2}\in\mathbb{R}^{c\times n}$ be output vectors (of the network) for $n$ samples. We can then approximate CKA by inserting $A=(A^{1/2})^{T}A^{1/2}$ and $B=(B^{1/2})^{T}B^{1/2}$ into the above equation. Because we already have the square root matrices, the computation cost of CKA is $\mathcal{O}(npc)$.
Recall that $n$ can be any number, and can be sampled from both train and test set. As $A$ and $B$ are at most rank $p$ and $c$ respectively, $n>p$ suffices in practice. Because $p\approx 10^{3}$ in many models, we can expect $\mathcal{O}(p^{2}c)$ computation. In an author’s laptop, it can be computed in less than a minute for all experiments. For NTK, $p$ is no longer the width of the last layer, but total number of parameters, quickly becoming infeasible.
E.2 Visualization
Time complexity.
The algorithm is computationally efficient in practice. The main cost in computing Eq.˜6 is the SVD, typically with time complexity $\mathcal{O}(p^{3})$. A good approximation requires only $n=\mathcal{O}(p)$ datapoints—more than the last layer width. Performing SVD on $\phi$ instead of $\phi\phi^{T}$ yields $U$ and $\sqrt{S}$ with $\mathcal{O}(p^{3})$ cost. Because last layers are typically narrow (e.g., $p\approx 10^{3}$) and the algorithm requires just $n=\mathcal{O}(p)$ forward passes, the overall computation is lightweight. In an author’s laptop, the computation for all experiments takes less than a few minutes.
Appendix F Examples of our visualization method
@@
Figure 11: Lazy/rich transition in target downscaling, illustrating how lazier dynamics (ii) use more features and (iii) show a slower decay of eigenvalues (ignoring $\rho_{1}$ of constant function, see Appendix˜D).
Figure 12: Lazy/rich transition in grokking. A 2-layer transformer is trained on the modular $p$ division task. The inset in the middle shows the training and test accuracies, where green and red vertical lines indicate before and after grokking (steps 200 and 3000). Our metric in square brackets shows the transition into the rich regime, while our visualization shows a clear difference in the use of features.
@@
Table˜4 shows that low-rank inductive bias is strong in vision tasks, where the number of significant features always equals $C$ (the number of orthogonal functions). See Eq.˜9 in Section˜B.4 for the $D_{eff}(\rho)$ expression. We also observe that our measure $\mathcal{D}_{LR}$ is small in these setups.
@@
Fig.˜13 shows that weight decay indeed aids low-rank representation and smaller $\mathcal{D}_{LR}$. However, the model without any weight decay already shows sufficiently low-rank representations, suggesting that dynamical inductive bias, not the weight decay is the main driving cause of low-rank representations.
@@
As studied in Saxe et al. (2022), the architecture and dataset pair can influence the performance even when dynamics are greedy (race toward shared representation). In Fig.˜14, we examine the dynamics of ResNet18 and a 4-layer MLP trained on CIFAR-10. In Fig.˜14(a), the training dynamics is restricted to the first 10 significant features, consistent with theoretical work in linear neural networks. In Fig.˜14(b), the training is mainly focused on the first 10 features, but we observe that more features are used as training progresses, more similar to the dynamics of linear models. We speculate that the lack of ability to feature learn in the earlier layers leads to the use of additional features and thus leads to lazier training dynamics.
Rich dynamics $\neq$ generalization.
In Fig.˜1, we showed an example where rich dynamics led to poorer performance. Fig.˜16 shows a similar experiment in which both rich and lazy models achieve near-perfect performance ($95\%$ test accuracy). Our visualization clearly shows the different usages of features even when the performances are similar.
Rich dynamics without underlying data structure.
Fig.˜15 illustrates that rich dynamics can occur independently of representation enhancement or underlying data structure. We trained ResNet18 on CIFAR-10 with varying levels of label shuffling and observed that even with fully randomized labels, the model enters the rich regime. This suggests that the dynamical low-rank bias is strong enough to collapse expressive networks into minimal representations, consistent with prior observations of neural collapse under random labeling (Zhu et al., 2021; Hui et al., 2022).
@@
Fig.˜17 shows that our method extends to models trained on cross-entropy loss, while the mathematical interpretation is less straightforward.
Figure 17: Training with cross-entropy loss. We apply our visualization method to a ResNet18 trained on CIFAR-10 with cross-entropy loss. Similar to the MSE-trained model, it achieves small value of $\mathcal{D}_{LR}$ and shows concentration of feature usage and significance See Fig.˜2(a) for comparison. The notable difference is in the speed difference among the first $C$ eigenfunctions. In MSE, the first $C$ eigenfunctions show nearly isotropic growth from epoch 1, while for CE, there is a notable difference in $\hat{\Pi}(k)$ and $\rho_{k}$ depending on $k$ until epoch 5. This may arise from the non-linearity coming from Cross entropy. Insets display all 512 features rather than just the top 25.
Appendix G Related works on visualization
Eigendecomposition has long been a standard tool in science and engineering (Buckley and Xu, 1990; Amadei et al., 1993; te Vrugt and Wittkowski, 2020), demonstrating its value for visualization. Building on this, related empirical works have appeared (Raghu et al., 2017; Morcos et al., 2018; Kornblith et al., 2019). For example, Singular vector canonical correlation analysis (Raghu et al., 2017) resembles eigendecomposition but focuses on performance along leading directions rather than on the number of features used to represent the learned function (i.e., lacking $\hat{\Pi}$). (Morcos et al., 2018; Kornblith et al., 2019) compare different networks, whereas our focus is on before- and after-last-layer activations. Importantly, these approaches were not developed as visualization tools for analyzing rich dynamics.
Similarly, visualizations related to neural collapse use eigendecomposition (Papyan, 2020), but are inherently class-dependent, whereas our metric and visualization are class-independent. See Section˜D.3 and Fig.˜10.
The closest line of work is kernel-based visualization (Bordelon et al., 2020; Canatar et al., 2021; Simon et al., 2023), especially (Canatar and Pehlevan, 2022). However, their visualization focuses on cumulative power and the eigenvalue spectrum, whereas ours additionally separates dynamics from performance, providing a complementary perspective. Additionally, the rich dynamics in the literature refers to change in NTK, not the collapse (low-rank bias) of feature kernel.
While visualization is part of our contribution, our main contribution lies in introducing $\mathcal{D}_{LR}$ as a metric of richness and presenting supporting empirical results. Visualization serves as a complementary tool to illustrate the dynamics and facilitate future studies.
Appendix H Empirical details
Here, we share information on our practical setups and their statistical significance. The code can be found at https://github.com/yoonsoonam119/PrEF.
H.1 Statistical significance
@@
Train acc. $\uparrow$ 100 $(\pm 0)$% 98.7 $(\pm 0.11)$ %
Test acc. $\uparrow$ 10.0 $(\pm 4.3\cdot 10^{-2})$% 74.4 $(\pm 0.19)$ %
$\mathcal{D}_{LR}$ $\downarrow$ $8.7\cdot 10^{-3}$ $(\pm 4.0\cdot 10^{-4})$ $6.3\cdot 10^{-1}$ $(\pm 3.2\cdot 10^{-3})$
Table 6: Statistical significance of Table 1
@@
[…]
$\bm{\mathcal{D}_{LR}}\downarrow$ 0.59 $(\pm 1.6\cdot 10^{-2})$ 1.0 $(\pm 6.4\cdot 10^{-5})$
$\mathcal{S}_{init}\downarrow$ 1.0 $(\pm 0)$ 0.20 $(\pm 1.1\cdot 10^{-2})$
$\|\theta\|_{F}^{2}\downarrow$ 3.1 $\cdot 10^{3}$ $(\pm 3.5\cdot 10^{2})$ 2.2 $\cdot 10^{-5}$ $(\pm 2.5\cdot 10^{-9})$
$NC1\downarrow$ $1.2\cdot 10^{5}$ $(\pm 7.2\cdot 10^{2})$ $7.5\cdot 10^{-14}$ $(\pm 1.5\cdot 10^{-14})$
$\mathrm{Tr}(\Sigma_{W})\downarrow$ $15.6$ $(\pm 1.4)$ $7.1\cdot 10^{-18}$ $(\pm 1.1\cdot 10^{-18})$
Table 7: Statistical significance of Table 2
@@
[…]
$\bm{\mathcal{D}_{LR}}\downarrow$ $\mathbf{4.9\cdot 10^{-2}}$ $(\pm 4.4\cdot 10^{-3})$ $\mathbf{1.1\cdot 10^{-1}}$ $(\pm 9.4\cdot 10^{-3})$ $\mathbf{5.6\cdot 10^{-1}}$ $(\pm 6.1\cdot 10^{-2})$
$\mathcal{S}_{init}\downarrow$ $6.8\cdot 10^{-2}$ $(\pm 2.1\cdot 10^{-3})$ $4.1\cdot 10^{-2}$ $(\pm 4.5\cdot 10^{-3})$ $5.2\cdot 10^{-2}$ $(\pm 1.8\cdot 10^{-2})$
$\|\theta\|_{F}^{2}\downarrow$ $3.4\cdot 10^{3}$ $(\pm 1.1\cdot 10^{1})$ $3.2\cdot 10^{3}$ $(\pm 4.4)$ $3.2\cdot 10^{3}$ $(\pm 2.0)$
$NC1\downarrow$ $2.3\cdot 10^{4}$ $(\pm 4.1\cdot 10^{3})$ $3.2\cdot 10^{3}$ $(\pm 4.3\cdot 10^{2})$ $8.1\cdot 10^{2}$ $(\pm 1.4\cdot 10^{2})$
$\mathrm{Tr}(\Sigma_{W})\downarrow$ $2.0$ $(\pm 0.25)$ $3.1\cdot 10^{-1}$ $(\pm 2.3\cdot 10^{-2})$ $1.2\cdot 10^{-1}$ $(\pm 1.2\cdot 10^{-2})$
Table 8: Statistical significance of Table 3
Task Architecture Condition $\mathcal{D}_{LR}$
Mod 97 2-layer transformer Step 200 (before grokking) 0.51 $(\pm 0.04)$
Step 3000 (after grokking) 0.11 $(\pm 0.01)$
CIFAR-100 ResNet18 learning rate = 0.005 0.053 $(\pm 0.01)$
learning rate = 0.05 0.025 $(\pm 0.004)$
learning rate = 0.2 0.039 $(\pm 0.007)$
CIFAR-10 ResNet18 weight decay = 0 0.05 $(\pm 0.003)$
weight decay = $10^{-4}$ 0.015 $(\pm 0.003)$
weight decay = $10^{-3}$ 0.003 $(\pm 0.001)$
CIFAR-10 ResNet18 . 0.026 $(\pm 0.003)$
MLP 0.48 $(\pm 0.07)$
CIFAR-10 ResNet18 no label shuffling 0.031 $(\pm 0.004)$
10% label shuffling 0.042 $(\pm 0.003)$
full label shuffling 0.034 $(\pm 0.004)$
MNIST CNN full backpropagation 0.043 $(\pm 0.005)$
last layer only training 0.51 $(\pm 0.008)$
CIFAR-100 VGG-16 without batch normalization 0.66 $(\pm 0.01)$
with batch normalization 0.073 $(\pm 0.002)$
H.2 Discussion on NC1 measure
In Tables˜6 and 7, the NC1 measure is numerically unstable due to the pseudo-inverse of $\Sigma_{b}$. For stability, we used $\Sigma_{b}^{\dagger}\approx(\Sigma_{b}+10^{-4}I)^{\dagger}$. In addition to NC1, we consider a similar measure $\mathrm{Tr}(\Sigma_{W})$ Papyan et al. (2020) for collapse of within class variability, which is also reported in Tables˜6 and 7. The measure $\mathrm{Tr}(\Sigma_{W})$ also shows a similar trend to NC1.
H.3 Dataset details
We use publicly available datasets, including MNIST Deng (2012) and CIFAR-10/100 Krizhevsky et al. (2009) from PyTorch Paszke et al. (2019), and mod $p$ division task from https://github.com/teddykoker/grokking.
H.4 Model details
Our model implementations are based on publicly available code assets, including VGG16 from PyTorch Paszke et al. (2019), ResNet18 from Nakkiran et al. Nakkiran et al. (2021), and the modular division task from https://github.com/teddykoker/grokking.
As described in the main text, we only trained our models on MSE loss where the targets are one-hot vectors (up to scaling constant). The constant $\alpha$ is used to scale the output $y\rightarrow y/\alpha$.